OpenStudy (anonymous):

Medal for Answer Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form. f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5

OpenStudy (anonymous):

@dan815 @dsood15 @desiswag @DoShKa_SyRiA

OpenStudy (anonymous):

@lupita1995 @love_jessika15

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1. f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

when focus and directrix are given use the deinition of the parabola or locus method.. let (x,y) be the point on the reqd parabola then distance of (x,y) from (-2,4) is sqrt ( (x+2)^2+(y-4)^2 ) and distance of (x,y) from y=6 is (y-6) units so by definition of parabola we have sqrt ( (x+2)^2+(y-4)^2 ) = (y-6) squaring both sides we have (x+2)^2+(y-4)^2 =(y-6)^2 and now just simplify